OpenStudy (anonymous):

Jill Ley took out a loan for $60,000 to pay for her child's education. The loan must be repaid at the end of eight years in one payment with interest of 6%. The total amount Jill has to pay back at the end of the loan is

OpenStudy (anonymous):

you can use I = prt to solve this: I = simple interest (what you want to solve) P = principle (what you start off with) in this case 60,000 r = rate (in this case 0.06 or 6% t = time (in years) in this case 8 so your answer would be 28800 + 60,000 = $88,800 your answer would be $88,800

OpenStudy (anonymous):

wait don't use that answer if it's not simple annual interest... if it isn't just calculate 6% of 60,000 or 1000.